OpenStudy (anonymous):

tell weather the system has one solution , infinitely, many solution or no solution 5x - 3y = 10 10x + 6y = 20

OpenStudy (mathstudent55):

Subtract the second equation from the first equation. If you get a false statement, there are no solutions.

OpenStudy (anonymous):

so the answer is one solution

OpenStudy (mathstudent55):

Did you subtract the equations? What did you get?

OpenStudy (anonymous):

get only 2 left

OpenStudy (mathstudent55):

9x + 8y = 15 9x + 8y = 30 --------------(subtract) 9x - 9x = 0; 8y - 8y = 0; 15 - 30 = -15 0 + 0 = -15 0 = -15 is a false statement, so there is no solution.

OpenStudy (anonymous):

can u help me with the top i change the question @mathstudent55

OpenStudy (anonymous):

It should be one solution 5x-3y=10 is practically same as 10x-6y=20 (when multiplied by 2 both sides). So when you solve the system of equations, 10x-6y=20 10x+6y=20 You can add the two equation to get x value, i.e. 20x = 40 Therefore x = 2 when x=2, you get y=0 so yeah.. one solution